# A Gentle Wakeup Call Symmetry Breaking with Less Collision Cost

### v2

@@ Abstract
Prior work on wakeup primarily analyzes latency—the number of slots until the first success. However, in many modern systems, each collision incurs a nontrivial delay, $\mathcal{C}$, which prior analyses neglect. Consequently, although existing algorithms achieve polylogarithmic-in-$n$ latency, they still suffer a delay of $\Omega(\mathcal{C})$ due to collisions.
@@
1 Introduction
In the wakeup problem, devices are either active (awake) or inactive (asleep). Each active device has a packet to send on a shared channel, while inactive devices are limited to listening. The first successful transmission serves as a “wakeup call”, awakening all inactive devices and resolving the problem [2].
@@ 1 Introduction
Why Should We Care About Collisions? The delay incurred by a collision can vary across different systems. As an example, consider WiFi (IEEE 802.11) networks; here, both sending a packet and the delay from a collision consumes time, say $\mathcal{C}$, proportional to the packet size [1]. In this setting, wakeup incurs an unavoidable delay of $\mathcal{C}$, since a single success is required. Yet, reducing the delay caused by collisions is still important. If this delay is, say, $X$, then throughput is $\mathcal{C}/(\mathcal{C}+X)$. So, if $X=\Theta(\mathcal{C})$, then throughput is sub-optimal by a constant factor. However, if say $X=\Theta(\sqrt{\mathcal{C}})$, then throughput is $\mathcal{C}/(\mathcal{C}+\Theta(\sqrt{\mathcal{C}}))\geq 1-\Theta(1/\sqrt{\mathcal{C}})$; that is, throughput approaches the optimal value of $1$ as $\mathcal{C}$ grows large. [^1] Empty slots will also degrade throughput, but in WiFi, these slots are negligible compared to the time of sending a packet or a collision. This illustrates how minimizing collisions improves performance.
More generally, over long-range links, where the propagation delays are large, a collision means devices can wait a long round-trip time before realizing the transmission has failed. Also, in intermittently-connected mobile wireless networks [21], failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range. Finally, in environments where there is some fluctuation in the delay (e.g., arising from fluctuating channel conditions or periods of heavy network usage), $\mathcal{C}$ can be interpreted as an upper bound on this delay that holds under most operating conditions.
1.1 Model and Notation
@@
No Collision Detection. There is no mechanism to distinguish between an empty slot and a collision—that is, no collision detection (no-CD). In many practical systems, collision detection (CD) is unavailable or unreliable. For example, in WiFi networks, received signal strength may offer a weak form of CD, but it is error-prone. Consequently, many prior results address the challenging no-CD setting (see the survey [2] and related discussion in Section 2).
@@
Metrics. We care about latency, which refers to the time between the first packet injection and the first success. Additionally, each collision incurs a known delay of $\mathcal{C}$ $\geq 4$. [^2] The value $4$ is chosen to simplify the analysis. We also note that only $\mathcal{C}=\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff [13, 10, 2]) for the standard model will be asymptotically optimal. To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree is of the polynomial is unknown. Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
@@ 1.2 Our Results
  2. 2.
    Else, the expected latency is $O(\log^{4+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{4+\frac{3}{2\epsilon})}n)$.
@@
2 Related Work
Wakeup, or leader election, addresses the fundamental problem of symmetry breaking. Early work by Willard [20] establishes a $\lg\lg n+O(1)$ upper bound with a matching lower bound for fair algorithms under collision detection (CD).
In contrast, without CD, a common theme is that the latency increases to logarithmic or polylogarithmic in $n$. For starters, without collision detection and when $n$ is known, Kushilevitz and Mansour [16, 15] prove a lower bound of $\Omega(\log n)$ expected latency. Newport [18, 19], in work that unifies many prior lower-bound results, shows $\Omega(\log n)$ expected latency and $\Omega(\log^{2}n)$ latency w.h.p. for general randomized algorithms. Complementing these lower bounds, Bar-Yehuda et al. [3] achieve a matching upper bound of $O(\log n)$ expected latency and w.h.p. $O(\log^{2}n)$ latency.
Farach-Colton et al. [7] obtain $\Omega(\log n\log(1/\varepsilon))$ latency for algorithms without CD, succeeding with probability at least $1-\varepsilon$. A complementary by Jurdziński and Stachowiak [14] demonstrates an algorithm that succeeds with probability at least $1-\varepsilon$ with $O(\log n\log(1/\varepsilon))$ latency, without CD, but assuming unique labels or knowledge of $n$.
With multiple channels and collision detection, Fineman et al. [9] give an efficient contention-resolution algorithm. In the SINR model [17], Fineman et al. [8] achieve logarithmic latency and establish matching lower bounds. Chang et al. [6] examine energy efficiency under varying feedback models, while Gilbert et al. [11] study the effect of system-size prediction, deriving corresponding bounds.
Research on the cost of collisions is still in its early stages. Anderton et al. [1] show that collisions create substantial discrepancies between theoretical and observed performance in backoff algorithms, underscoring the need to account for collision costs. Building on this, Biswas et al. [4] provide one of the first analyses, addressing the problem of contention resolution and deriving bounds on expected latency and total expected collision cost.
@@ 3 Algorithm Specification
Doubling phase. At the beginning of this phase, $w_{\mbox{\tiny cur}}$ starts at $2w_{0}$. Each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of a sample (Line 1 and 1). If a success occurs, the execution terminates; otherwise, the window doubles and another sample is executed. This process continues until $2^{j}$ samples (each corresponding to a doubling of the window) are performed, where $j=1$ initially, and increments at the end of each iteration.
3.1 Context and Overview
@@
3.1.1 A Baseline via Backoff.
A natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$. At $i=\lg(n)$, the probability of a success in a slot becomes constant. An example of its use in solving wakeup is given by Jurdziński and Stachowiak [14], who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours (without a notion of collision cost); however, each active packet starts with a high (i.e., constant) sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
@@
3.1.2 Design Choices and Overview.
@@
Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as the size of a good window. Informally, once packets reach a good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$. Therefore, over $\Theta(\sqrt{\mathcal{C}})$ slots, we expect a success. At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots. This observation dictates the size of our samples, which are $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ and appear in lines 1 and 1 in the halving phase, and lines 1 and 1 in the doubling phase of our pseudocode.
@@
Why Iterations? Although AH is likely to terminate when a good window is reached, there is some small probability of failure. Suppose we are in Case 1, but this failure occurs in the halving phase. Then, this small probability of failure yields good expected total cost of collisions as the window size decreases past the good window down to $2$ (see Lemma 3). However, if the halving phase is never revisited, the packets may accrue intolerable latency by executing the doubling phase without success (as the sending probability drops off exponentially fast). Conversely, suppose we are in Case 2 and this failure occurs. Then, if we never reset $w_{\mbox{\tiny cur}}$ to (roughly) $w_{0}$, the latency costs may again be high. Therefore, iterations facilitate good expected latency (see Lemma 6).
Parameter $\epsilon$. Throughout, we treat $\epsilon$ as a small constant for the purposes of our analysis. Yet, it is worth pointing out how it interacts with the latency and the expected total cost of collisions. The former is improved by making $\epsilon$ small, since this shortens the halving phase. So, if aiming high pays off, then a smaller $\epsilon$ can reduce latency. However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, degree of the polynomial increases with the reciprocal of $\epsilon$.
4 Upper Bound
In this section, we analyze the performance of AH. Define $\eta$ $\max\{n,\mathcal{C}\}$. A good window was loosely discussed in Section 3.1.2; we define it as any window with size in $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$. Throughout, we assume that $n$ is sufficiently large. We make use of the following well-known facts.
Fact 1
The following inequalities hold.
  1. (a)
    For any $x$, $1-x\leq e^{-x}$,
  2. (b)
    For any $0\leq x\leq 1/2$, $1-x\geq e^{-2x}$.
We also make use of the following standard Chernoff bounds.
Theorem 4.1
(Goemans [12]) Let $X=\sum_{i}X_{i}$, where $X_{i}$ are Bernoulli indicator random variables with probability $p$. The following holds:
$Pr\left(X\geq(1+\delta)E[X]\right)$ $\leq\exp\left(-\frac{\delta^{2}E[X]}{2+\delta}\right)\mbox{~~~for any $\delta>0$}$
and
[…]
The next lemma argues that success is likely in a good window.
Lemma 1
@@ 4 Upper Bound
Proof
Let $w^{*}\in[n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$ (recall Section 3). The probability of a success in any fixed slot of this good window is at least:
[…]
$\geq\left(\frac{1}{3e\sqrt{\mathcal{C}}}\right)\mbox{~~~since $\mathcal{C}\geq 4$.}$
Let $X_{i}=1$ if the $i$-th slot of the sample contains a success; otherwise, $X_{i}=0$. The expected number of successes in the sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots is at least:
$E[X]\geq\left(\frac{d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})}{3e\sqrt{\mathcal{C}}}\right)$ $=\left(\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$
where $d^{\prime}$ is a constant that can be made arbitrarily large depending on $d$.
We then employ Theorem 4.1, with $\delta=1/2$, to obtain:
$Pr\left(X\leq(1-\delta)\frac{d\ln(w_{\mbox{\tiny cur}})}{3e}\right)$ $\leq Pr\left(X\leq\frac{d\ln(w_{\mbox{\tiny cur}})}{6e}\right)$
$\leq e^{-\frac{{(1/2)}^{2}d\ln(w_{\mbox{\tiny cur}})}{(5/2)3e}}$
$=e^{-\frac{d\ln(n\sqrt{\mathcal{C}})}{30e}}$
$\leq(n\sqrt{\mathcal{C}})^{-d^{\prime}}$
$\leq\eta^{-d^{\prime\prime}}$
where the third line follows from $w_{\mbox{\tiny cur}}=\Theta(\ln n\sqrt{\mathcal{C}})$, and $d^{\prime}$ and $d^{\prime\prime}$ can be arbitrarily large constants for $d$ is sufficiently large.
4.1 Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
@@
Lemma 2
(Biswas et al. [4]) Consider any slot $t$ where there exist $m$ active packets, each sending with the same probability $p$, where $p<1/m$. The probability of a collision in slot $t$ is at most $\frac{2m^{2}p^{2}}{\left(1-mp\right)}$.
Lemma 3
@@
$G+\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ (1)
Upper Bounding $B_{j}$. Clearly, $B_{j}$ is upper-bounded by the cost of having a collision in every single slot of iterations $1$ through $j$. In any iteration $k$, there is a sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots for each of the at most $\ln(w_{0})$ windows; thus, over iterations $1$ through $j$, the halving phases contributes $O(j\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0}))$ slots. Additionally, in iteration $k$, the corresponding doubling phase executes $2^{k}$ samples, each consisting of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots; thus, by the sum of a geometric series, over iterations $1$ through $j$, the doubling phases contributing $O(\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j})$ slots. By the above:
$B_{j}$ $=O\left(j\,d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})+\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})2^{j}\right)\cdot\mathcal{C}$
$=O\left(j\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\log(w_{0})2^{j}\right)\cdot\mathcal{C}$
$=O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
since $\log(w_{0})=\log(2^{\mathcal{C}^{\epsilon}})=O(\mathcal{C}^{\epsilon})$, and $\ln(w_{\mbox{\tiny cur}})\leq\ln(2^{j}w_{0})=O(j\mathcal{C}^{\epsilon})$.
By Lemma 1, the $q=1/\eta^{c}$, where recall that $\eta=\max\{n,\mathcal{C}\}$ where we can make the constant $c>0$ as large as we wish. Thus,
$\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ $\leq\sum_{j=1}^{\infty}(1/\eta^{c})^{j}\cdot O(C^{(3/2)+2\epsilon}\,j^{2}\,2^{j})$
@@
Upper Bounding $G$. To upper bound $G$, for each slot $i$ in the first halving phase of the first iteration, let $X_{i}$ be an independent random variable where $X_{i}=1$ if the $i$-th slot contains a collision; otherwise, $X_{i}=0$. Recall that the windows of this first halving phase are repeatedly halved. For a slot in a window of size $w=w_{0}/2^{k}$, for a positive integer $k\geq 0$, by Lemma 2, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$), we have:
$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}$
$=\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}$
@@
Note that the number of windows executed starting from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to size $n\sqrt{\mathcal{C}}$, is at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$; for ease of presentation, let $x=C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$. By the above, we can derive:
@@
$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt\frac{2^{2k}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$ (2)
$\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt2^{2k}$ (3)
$=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$ (4)
$<\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{8}{\mathcal{C}}$
@@
In the above, line 2 holds since $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$, since we are in Case 1. Line 3 holds for any $\mathcal{C}\geq 4$, and line 4 holds by the sum of the geometric series. Thus, Equation 1 is bounded as claimed.
The next lemma bounds the expected latency of our algorithm, and its argument is similar to that of Lemma 3, although an extra $O(\mathcal{C}^{\epsilon})$-factor appears. Informally, this occurs since, in counting slots for the first halving phase, we have a sample size of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})=d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $\leq d\sqrt{\mathcal{C}}\ln(w_{0})$ $d\sqrt{\mathcal{C}}C^{\epsilon}=O(\mathcal{C}^{(1/2)+\epsilon})$, and we have $\log_{2}w_{0}=O(C^{\epsilon})$ samples.
Lemma 4
@@
4.2 Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
@@
$>\mathcal{C}^{\epsilon}-(\mathcal{C}^{\epsilon}/2).$ (5)
To justify the second line, let $\delta\in(0,1)$, and note that:
$(\lg\mathcal{C})/2$ $<\mathcal{C}^{\epsilon}/2$
@@
so long as $\epsilon>\frac{\lg\lg(\mathcal{C})}{\lg\mathcal{C}}$, which holds by assumption since $\epsilon$ is a constant, and this is well-defined since $\mathcal{C}\geq 4$. Returning to Equation 5, we have that:
$(2\lg n)^{1/\epsilon}$ $>\mathcal{C}.$
as desired, since $\epsilon$ is a constant.

### v3

@@ Abstract.
Prior work primarily analyzes latency—the number of slots until the first success. However, in many modern systems, each collision incurs a nontrivial delay, $\mathcal{C}$, which prior analyses neglect. Consequently, although existing algorithms achieve polylogarithmic-in-$n$ latency, they still suffer a delay of $\Omega(\mathcal{C})$ due to collisions.
@@
1. Introduction
In the wakeup problem, devices are either active (awake) or inactive (asleep). Each active device has a packet to send on a shared channel, while inactive devices are limited to listening. The first successful transmission serves as a “wakeup call”, awakening all inactive devices and resolving the problem (Banicescu et al., 2024).
@@ 1. Introduction
Why Should We Care About Collisions? The delay incurred by a collision can vary across different systems. As an example, consider WiFi (IEEE 802.11) networks; here, both sending a packet and the delay from a collision consume time, say some large value $\mathcal{C}$, proportional to the packet size (Anderton et al., 2021). In this setting, wakeup incurs an unavoidable delay of $\mathcal{C}$, since a single success is required. Yet, reducing the delay caused by collisions is still important. If this delay is, say, $X$, then $\mathcal{C}/(\mathcal{C}+X)$ provides a measure of how much time is wasted due to collisions. So, if $X=\Theta(\mathcal{C})$, then this ratio is suboptimal by a constant factor. However, if $X=\Theta(\sqrt{\mathcal{C}})$, then $\mathcal{C}/(\mathcal{C}+\Theta(\sqrt{\mathcal{C}}))\geq 1-\Theta(1/\sqrt{\mathcal{C}})$; that is, the ratio approaches the optimal value of $1$ as $\mathcal{C}$ grows large.
More generally, over long-range links, where the propagation delays are large, a collision means devices can wait a long round-trip time before realizing the transmission has failed. Also, in intermittently-connected mobile wireless networks (Zhang, 2006), failed communication due to a collision may result in the sender having to wait until the intended receiver is again within transmission range. Finally, in environments where there is some fluctuation in the delay (e.g., arising from fluctuating channel conditions or periods of heavy network usage), $\mathcal{C}$ can be interpreted as an upper bound on this delay that holds under most operating conditions.
1.1. Model and Notation
@@
No Collision Detection. There is no mechanism to distinguish between an empty slot and a collision—that is, no collision detection (no-CD). In many practical systems, collision detection (CD) is unavailable or unreliable. For example, in WiFi networks, received signal strength may offer a weak form of CD, but it is error-prone. Consequently, many prior results address the challenging no-CD setting (see the survey (Banicescu et al., 2024) and related discussion in Section 2).
@@
Metrics. We care about latency, which refers to the time between the first packet injection and the first success. Additionally, each collision incurs a known delay of $\mathcal{C}$ $\geq 4$. [^1] The value $4$ is chosen to simplify the analysis. We also note that only $\mathcal{C}=\omega(1)$ is interesting; otherwise, an optimal algorithm (such as sawtooth backoff (Greenberg and Leiserson, 1985; Geréb-Graus and Tsantilas, 1992; Banicescu et al., 2024)) for the standard model will be asymptotically optimal. To capture the worst-case performance, $\mathcal{C}$ is chosen by the adversary with knowledge of $n$, subject to $\mathcal{C}=O({\texttt{poly}({n})})$, where the degree of the polynomial is unknown. Note that this reveals no information about $n$, whose value is unknown to the packets a priori.
@@ 1.2. Our Results
  2. (2)
    Else, the expected latency is $O(\log^{5+\frac{1}{2\epsilon}}n)$ and the expected total cost of collisions is $O(\log^{5+\frac{3}{2\epsilon}}n)$.
@@
2. Related Work
Wakeup, or leader election, addresses the fundamental problem of symmetry breaking. Early work by Willard (Willard, 1986) establishes a $\lg\lg n+O(1)$ upper bound with a matching lower bound for fair algorithms under collision detection (CD).
In contrast, without CD, a common theme is that the latency increases to logarithmic or polylogarithmic in $n$. For starters, without collision detection and when $n$ is known, Kushilevitz and Mansour (Kushilevitz and Mansour, 1998, 1993) prove a lower bound of $\Omega(\log n)$ expected latency. Newport (Newport, 2014a, b), in work that unifies many prior lower-bound results, shows $\Omega(\log n)$ expected latency and $\Omega(\log^{2}n)$ latency w.h.p. for general randomized algorithms. Complementing these lower bounds, Bar-Yehuda et al. (Bar-Yehuda et al., 1992) achieve a matching upper bound of $O(\log n)$ expected latency and w.h.p. $O(\log^{2}n)$ latency.
Farach-Colton et al. (Farach-Colton et al., 2006) obtain $\Omega(\log n\log(1/\varepsilon))$ latency for algorithms without CD, succeeding with probability at least $1-\varepsilon$. A complementary by Jurdziński and Stachowiak (Jurdziński and Stachowiak, 2002) demonstrates an algorithm that succeeds with probability at least $1-\varepsilon$ with $O(\log n\log(1/\varepsilon))$ latency, without CD, but assuming unique labels or knowledge of $n$.
With multiple channels and collision detection, Fineman et al. (Fineman et al., 2016b) give an efficient contention-resolution algorithm. In the SINR model (Moscibroda and Wattenhofer, 2006), Fineman et al. (Fineman et al., 2016a) achieve logarithmic latency and establish matching lower bounds. Chang et al. (Chang et al., 2019) examine energy efficiency under varying feedback models, while Gilbert et al. (Gilbert et al., 2021) study the effect of system-size prediction.
Research on the cost of collisions is still in its early stages. Anderton et al. (Anderton et al., 2021) show that collisions create substantial discrepancies between theoretical and observed performance in backoff algorithms, underscoring the need to account for collision costs. Building on this, Biswas et al. (Biswas et al., 2024a) provide one of the first analyses, addressing the problem of contention resolution and deriving bounds on expected latency and total expected collision cost.
@@ 3. Algorithm Specification
Doubling phase. At the beginning of this phase, $w_{\mbox{\tiny cur}}$ starts at $2w_{0}$. Each packet sends with probability $\Theta(1/w_{\mbox{\tiny cur}})$ in each slot of a sample (Line 1 and 1). If a success occurs, the execution terminates; otherwise, the window doubles and another sample is executed. This process continues until $j$ samples (each corresponding to a doubling of the window) are performed, where $j=1$ initially, and increments at the end of each iteration.
3.1. Context and Overview
@@
3.1.1. A Baseline via Backoff.
A natural approach for breaking symmetry is to “backoff”, where each packet sends with probability $1/2^{i}$ in slot $i\geq 0$. At $i=\lg(n)$, the probability of a success in a slot becomes constant. An example of its use in solving wakeup is given by Jurdziński and Stachowiak (Jurdziński and Stachowiak, 2002), who achieve $O(n/\log n)$ latency with bounded error $\varepsilon$ in a similar model to ours (without a notion of collision cost); however, each active packet starts with a high (i.e., constant) sending probability, which inflicts $\Omega(\mathcal{C})$ cost.
@@
3.1.2. Design Choices and Overview.
@@
Consider a window of size $w^{*}=\Theta(n\sqrt{\mathcal{C}})$; we refer to this as the size of a good window. Informally, once packets reach a good window, the chance of a success in any slot of the window is at least $\binom{n}{1}(1/w^{*})\approx 1/\sqrt{\mathcal{C}}$. Therefore, over $\Theta(\sqrt{\mathcal{C}})$ slots, we expect a success. At the same time, the probability of a collision is approximately $\binom{n}{2}(1/w^{*})^{2}\approx 1/\mathcal{C}$, and thus the expected collision cost per slot is $1$, and $O(\sqrt{\mathcal{C}})$ over $\Theta(\sqrt{\mathcal{C}})$ slots. This observation dictates the size of our samples.
@@
Why Iterations? Although AH is likely to terminate when a good window is reached, there is some small probability of failure. Suppose we are in Case 1, but this failure occurs in the halving phase. Then, this small probability of failure yields good expected total cost of collisions as the window size decreases past the good window down to $2$ (see Lemma 4.3). However, if the halving phase is never revisited, the packets may accrue intolerable latency by executing the doubling phase without success (as the sending probability drops off exponentially fast). Conversely, suppose we are in Case 2 and this failure occurs. Then, if we never reset $w_{\mbox{\tiny cur}}$ to (roughly) $w_{0}$, the latency costs may again be high. Therefore, iterations facilitate good expected latency (see Lemma 4.6).
Parameter $\epsilon$. Throughout, we treat $\epsilon$ as a small constant for the purposes of our analysis. Yet, it is worth pointing out how it interacts with the latency and the expected total cost of collisions. The former is improved by making $\epsilon$ small, since this shortens the halving phase. So, if aiming high pays off, then a smaller $\epsilon$ can reduce latency. However, we pay for this in terms of the latter metric; specifically, when $n\sqrt{\mathcal{C}}>w_{0}$ and the algorithm looks to succeed in the doubling phase, we have $\mathcal{C}=O(\log^{1/\epsilon}n)$. Thus, the degree of the polynomial increases with the reciprocal of $\epsilon$.
4. Upper Bound
In this section, we analyze the performance of AH; due to space constraints, we omit several proofs. Define $\eta$ $\max\{n,\mathcal{C}\}$. A good window was loosely discussed in Section 3.1.2; we define it as any window with size in $n\sqrt{\mathcal{C}},3n\sqrt{\mathcal{C}}]$. The next lemma argues that success is likely in a good window.
Lemma 4.1.
@@ 4. Upper Bound
4.1. Case 1: $n\sqrt{\mathcal{C}}\leq w_{0}$.
@@
Lemma 4.2.
(Biswas et al. (Biswas et al., 2024a)) Consider any slot $t$ where there exist $m$ active packets, each sending with the same probability $p$, where $p<1/m$. The probability of a collision in slot $t$ is at most $\frac{2m^{2}p^{2}}{\left(1-mp\right)}$.
Lemma 4.3.
@@
(1) $G+\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$
Upper Bounding $B_{j}$. Clearly, $B_{j}$ is upper-bounded by the cost of having a collision in every single slot of iterations $1$ through $j$. In any iteration $k$, there is a sample of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots for each of the at most $\lg(w_{0})$ windows; thus, over iterations $1$ through $j$, the halving phases contribute $O(j\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0}))$ slots. Additionally, in iteration $k$, the corresponding doubling phase executes $k$ samples, each consisting of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ slots; thus, the doubling phases contribute $O(j^{2}\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}}))$ slots. By the above:
$B_{j}$ $=O\left(j\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})+j^{2}\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\right)\cdot\mathcal{C}$
$=O\left(j^{2}\,\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})\lg(w_{0})\right)\cdot\mathcal{C}$
$=O\left(j^{3}C^{(3/2)+2\epsilon}\right)$
since $\log(w_{0})=\log(2^{\mathcal{C}^{\epsilon}})=O(\mathcal{C}^{\epsilon})$, and $\ln(w_{\mbox{\tiny cur}})\leq\ln(jw_{0})=O(j\mathcal{C}^{\epsilon})$.
By Lemma 4.1, $q=1/\eta^{c}$, where recall that $\eta=\max\{n,\mathcal{C}\}$ and we can make the constant $c>0$ as large as we wish. Thus,
$\sum_{j=1}^{\infty}q^{j}\cdot B_{j}$ $\leq\sum_{j=1}^{\infty}(1/\eta^{c})^{j}\cdot O(j^{3}C^{(3/2)+2\epsilon})$
@@
Upper Bounding $G$. To upper bound $G$, for each slot $i$ in the first halving phase of the first iteration, let $X_{i}$ be an independent random variable where $X_{i}=1$ if the $i$-th slot contains a collision; otherwise, $X_{i}=0$. Recall that the windows of this first halving phase are repeatedly halved. In a window of size $w=w_{0}/2^{k}$ where $k\geq 0$ is a positive integer, by Lemma 4.2, with $m=n$ and $p=1/w$ (since $w\geq n\sqrt{\mathcal{C}}$ and $\mathcal{C}\geq 4$, we satisfy the requirement that $p<1/m$):
$Pr(X_{i}=1)$ $\leq\frac{2(n/w)^{2}}{\left(1-(n/w)\right)}=\frac{2(n/(w_{0}/2^{k}))^{2}}{\left(1-(n/(w_{0}/2^{k}))\right)}.$
@@
Note that the number of windows executed starting from an initial size $w_{0}=2^{\mathcal{C}^{\epsilon}}$ down to size $n\sqrt{\mathcal{C}}$, is at most $C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$; for ease of presentation, let $x=C^{\epsilon}-\lg{n}-(\lg\mathcal{C})/2$. We have:
@@
(2) $\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{2n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt\frac{2^{2k}}{\left(1-\frac{1}{\sqrt{{\mathcal{C}}}}\right)}$
(3) $\leq\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\sum_{k=0}^{x}\hskip 0.0pt2^{2k}$
(4) $=\left(d\mathcal{C}^{(3/2)+\epsilon}\right)\frac{4n^{2}}{({2^{\mathcal{C}^{\epsilon}}})^{2}}\frac{4^{\mathcal{C}^{\epsilon}+1}}{(4-1)4^{\lg{n}}4^{\frac{\lg{\mathcal{C}}}{2}}}$
@@
In the above, line 2 holds since $n\sqrt{\mathcal{C}}\leq 2^{\mathcal{C}^{\epsilon}}$, since we are in Case 1. Line 3 holds for any $\mathcal{C}\geq 4$, and line 4 holds by the sum of the geometric series. Thus, Equation 1 is bounded as claimed. ∎
The next lemma bounds the expected latency of our algorithm, and its argument is similar to that of Lemma 4.3, although an extra $O(\mathcal{C}^{\epsilon})$-factor appears. Informally, this occurs since, in counting slots for the first halving phase, we have a sample size of $d\sqrt{\mathcal{C}}\ln(w_{\mbox{\tiny cur}})$ $\leq d\sqrt{\mathcal{C}}\ln(w_{0})$ $d\sqrt{\mathcal{C}}C^{\epsilon}=O(\mathcal{C}^{(1/2)+\epsilon})$, and we have $\log_{2}w_{0}=O(C^{\epsilon})$ samples.
Lemma 4.4.
@@
4.2. Case 2: $n\sqrt{\mathcal{C}}>w_{0}$.
@@
(5) $>\mathcal{C}^{\epsilon}-(\mathcal{C}^{\epsilon}/2).$
@@
where the second line holds so long as $\epsilon>\frac{\lg\lg(\mathcal{C})}{\lg\mathcal{C}}$, which is true by assumption since $\epsilon$ is a constant, and this is well-defined since $\mathcal{C}\geq 4$. Returning to Equation 5, we have that $(2\lg n)^{1/\epsilon}>\mathcal{C}$, as desired, since $\epsilon$ is a constant. ∎